# Holographic duality from Howe duality: Chern–Simons gravity as an ensemble of code CFTs

### v1

@@ Appendix D Higher-genus modular transformations and MacWilliams identities
Just as the genus-1 enumerator polynomials of even self-dual codes are invariant under the $S$ and $T$ transformations (47) and (48), which correspond to the genus-$1$ modular group, the genus-$g$ enumerator polynomial are invariant under transformations of the genus-$g$ modular group $Sp(2g,{\mathbb{Z}})$. Here we give explicit form of the generators of $\mathrm{Sp}(2g,{\mathbb{F}}_{p})$ acting on genus-$g$ codeword blocks (Chern-Simons wavefunctions), or equivalently, variables entering enumerator polynomials. All relations here in fact generalize to groups defined over ${\mathbb{Z}}_{k}$ for general $k$.
@@
$\Psi_{c_{1}\dots c_{g}}(\Omega,\xi,\bar{\xi})=\frac{1}{\Phi(\Omega)}\sum_{v_{I}\in\Lambda_{c_{I}}}e^{i\pi\,v_{I}^{T}\Pi_{IJ}v_{J}+2\pi i(p^{I}_{L}\cdot\xi^{I}-p^{I}_{R}\cdot\bar{\xi}^{I})+\pi\,\Xi_{IJ}(\xi^{T}_{I}\xi_{J}+\bar{\xi}^{T}_{I}\bar{\xi}_{J})},$ (188)
@@
$\Psi_{c_{1}\dots c_{g}}(\Omega^{\prime},\xi^{\prime},\bar{\xi}^{\prime})$ $=$ $\Psi_{c_{1}\dots c_{g}}(\Omega,\xi,\bar{\xi})\,e^{\frac{i\pi}{p}(c_{I},c_{J})T_{IJ}}.$ (191)
Corresponding $Sp(2g,{\mathbb{F}}_{p})$ group element are $\left(\begin{smallmatrix}1&T\\ 0&1\end{smallmatrix}\right)$.
@@
$\Omega\rightarrow(\Omega^{\prime})_{ij}$ =$=\left\{\begin{array}[]{cr}\Omega_{IJ}-\Omega_{IL}\Omega_{LJ}/\Omega_{LL},&I,J\neq L,\\ \Omega_{IL}/\Omega_{LL},&J=L,\\ \Omega_{LJ}/\Omega_{LL},&I=L,\\ -1/\Omega_{LL},&I=J=L,\end{array}\right.$ (196)
$\xi\rightarrow\xi^{\prime}_{i}$ =$=\left\{\begin{array}[]{cr}\xi_{I}-\xi_{L}\Omega_{LI}/\Omega_{LL},&I\neq L,\\ \xi_{L}/\Omega_{LL},&I=L,\end{array}\right.$ (199)
$\bar{\xi}\rightarrow\bar{\xi}^{\prime}_{i}$ =$=\left\{\begin{array}[]{cr}\bar{\xi}_{I}-\bar{\xi}_{L}{\bar{\Omega}}_{LI}/\bar{\Omega}_{LL},&I\neq L,\\ \bar{\xi}_{L}/\bar{\Omega}_{LL},&I=L,\end{array}\right.$ (202)
@@
This transformation corresponds to $Sp(2g)$ matrix
@@ Appendix E Explicit examples
In this appendix, we present additional examples that generalise the simplest non-trivial case of $[png]=[311]$ of section 3.3. We begin by presenting some commands in GAP GAP that are useful for working out the examples. Then we give complete examples of the theta correspondence for the cases $[png]=[312],[511],[321]$, and partial results for the case $[png]=[322]$.
@@ E.1 Method and manipulations in GAP
First we explain how to perform some of the computations behind the examples. In summary, we need to exhibit a finite-dimensional reducible representation of the group $G\times H$ on the space $\mathcal{X}$ of dimension $p^{2ng}$ whose basis are the vectors $\psi_{c_{1},\ldots,c_{g}}$. We take $G=O(n,n,{\mathbb{F}}_{p})=O^{+}(2n,{\mathbb{F}}_{p})$ (the meaning of $+$ was explained in appendix B) and $H=Sp(2g,{\mathbb{F}}_{p})$. These groups can be generated in GAP by specifying the quadratic/bilinear form. Here we give the commands for $[png]=[311]$. The orthogonal group and its character table are generated by the following commands: [^36] In GAP, double semicolon suppresses the output. We found that for the case $p=5$, $n=2$, $g=1$ there are some issues in executing the code given here. Note also that in GAP, as in ATLAS Atlas1985 , the orthogonal group is denoted “GO” (general orthogonal).
@@ E.1.1 Example at genus $g=2$
The space $\mathcal{X}_{+}^{[312]}$ has dimension $813^{4}=$ 81, [^39] Comparing with (169) ($N=4$), it is a direct sum of two irreps of the larger symplectic group $Sp(8,{\mathbb{F}}_{3})$ with dimensions $40$ (charge-conjugation odd) and $41$ (charge-conjugation even) respectively. and following the steps above, we find that under the product group $O^{+}(2,{\mathbb{F}}_{3})\times Sp(4,{\mathbb{F}}_{3})$ it has a decomposition,
@@
$Sp_{4,3}$ $[1]$ $[1]^{\prime}$ $[90]$ $[4320]$ $[6480]$ $[40]$cdots
$r_{1}$ $1$ $1$ $1$ $1$ $1$ $1$
$r_{4}$ $4$ $-4$ 00 00 00 $2\zeta-\zeta^{2}$
$r_{4}^{\prime}$ $4$ $-4$ 00 00 00 $2\zeta^{2}-\zeta$
$r_{5}$ $5$ $5$ $-3$ $1$ $-1$ $\zeta^{2}-2\zeta$
$r_{5}^{\prime}$ $5$ $5$ $-3$ $1$ $-1$ $\zeta-2\zeta^{2}$
$r_{6}$ $6$ $6$ $-2$ $-1$ 00 $-3$
$r_{10}$ $10$ $10$ $2$ $1$ 00 $2\zeta^{2}+5\zeta$
$r_{10}^{\prime}$ $10$ $10$ $2$ $1$ 00 $2\zeta+5\zeta^{2}$
$r_{15}$ $15$ $15$ $-1$ $-1$ $-1$ $6$
$r_{15}^{\prime}$ $15$ $15$ $7$ 00 $1$ $-3$
$r_{20}$ $20$ $-20$ 00 00 00 $-7$
$r_{20}^{\prime}$ $20$ $20$ $4$ $1$ 00 $2$
$r_{20}^{\prime\prime}$ $20$ $-20$ 00 00 00 $-5\zeta-8\zeta^{2}$
$r_{20}^{\prime\prime\prime}$ $20$ $-20$ 00 00 00 $-5\zeta^{2}-8\zeta$
$r_{20}^{\prime\prime\prime\prime}$ $20$ $-20$ 00 00 00 $\zeta-5\zeta^{2}$
$r_{20}^{\prime\prime\prime\prime\prime}$ $20$ $-20$ 00 00 00 $\zeta^{2}-5\zeta$
$r_{24}$ $24$ $24$ $8$ 00 00 $6$
$r_{30}$ $30$ $30$ $-10$ $-1$ 00 $3$
@@
Table 2: Truncated character table for the group $Sp(4,{\mathbb{F}}_{3})$ with $51840$ elements, displaying the two conjugacy classes with one element ($[1]$ and $[1]^{\prime}$), three additional conjugacy classes whose size is unique among the conjugacy classes, and one additional conjugacy class necessary to distinguish the representations $r_{20}^{\prime\prime}$ and $r_{20}^{\prime\prime\prime}$. The whole character table contains $34$ conjugacy classes. $\zeta=e^{2\pi i/3}$.
@@
$\ \left(\begin{smallmatrix}1&1&0&1\\ 2&2&2&0\\ 0&2&1&2\\ 1&0&1&2\end{smallmatrix}\right)\in[90],\ \left(\begin{smallmatrix}2&2&2&2\\ 2&2&0&1\\ 0&2&0&1\\ 0&1&2&2\end{smallmatrix}\right)\in[4320],\ \left(\begin{smallmatrix}1&1&2&0\\ 0&0&1&2\\ 2&2&1&1\\ 2&0&0&1\end{smallmatrix}\right)\in[6480],\ \left(\begin{smallmatrix}2&1&2&1\\ 2&0&1&2\\ 1&1&0&1\\ 1&1&2&2\end{smallmatrix}\right)\in[40].$ (210)
@@ E.2 Example for prime $p=5$
where the last term constitutes the charge-conjugation odd subspace of dimension $12$. Again, we see that $\pi_{1}\otimes r_{1}$ only appears once. The identification with bipartitions for unipotent irreps is the same as in the example $p=3$ in the main text, where now the dimension formula (118) and shows that $\dim r_{\emptyset,[1]}=5$.
$O_{2,5}^{+}$ $[1]$ $[2]$ $[2]^{\prime}$ $[2]^{\prime\prime}$ $[1]^{\prime}$
$\pi_{1}$ $1$ $1$ $1$ $1$ $1$
$\pi_{1}^{\prime}$ $1$ $-1$ $-1$ $1$ $1$
$\pi_{1}^{\prime\prime}$ $1$ $-1$ $1$ $-1$ $1$
$\pi_{1}^{\prime\prime\prime}$ $1$ $1$ $-1$ $-1$ $1$
$\pi_{2}$ $2$ 00 00 00 $-2$
@@
$SL_{2,5}$ $[1]$ $[12]$ $[12]^{\prime}$ $[1]^{\prime}$ $[12]^{\prime\prime}$ $[12]^{\prime\prime\prime}$ $[20]$ $[20]^{\prime}$ $[30]$
@@
$r_{2}$ 2 $-\xi^{4}-\xi$ $-\xi^{3}-\xi^{2}$ $-2$ $\xi^{4}+\xi$ $\xi^{3}+\xi^{2}$ $-1$ $1$ 00
$r_{2^{\prime}}$ $2$ $-\xi^{3}-\xi^{2}$ $-\xi^{4}-\xi$ $-2$ $\xi^{3}+\xi^{2}$ $\xi^{4}+\xi$ $-1$ $1$ 00
$r_{3}$ $3$ $-\xi^{3}-\xi^{2}$ $-\xi^{4}-\xi$ $3$ $-\xi^{3}-\xi^{2}$ $-\xi^{4}-\xi$ 00 00 $-1$
$r_{3^{\prime}}$ $3$ $-\xi^{4}-\xi$ $-\xi^{3}-\xi^{2}$ $3$ $-\xi^{4}-\xi$ $-\xi^{3}-\xi^{2}$ 00 00 $-1$
$r_{4}$ $4$ $-1$ $-1$ $4$ $-1$ $-1$ $1$ $1$ 00
$r_{4}^{\prime}$ $4$ $1$ $1$ $-4$ $-1$ $-1$ $1$ $-1$ 00
$r_{5}$ $5$ 00 00 $5$ 00 00 $-1$ $-1$ $1$
$r_{6}$ $6$ $-1$ $-1$ $-6$ $1$ $1$ 00 00 00
@@ E.3 Example at length $n=2$
$O_{4,3}^{+}$ $[1]$ $[1]^{\prime}$ $[18]$ $[144]$ $[72]$ $[72]^{\prime}$ $[72]^{\prime\prime}$cdots
$\pi_{1}$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
@@
$\pi_{2}$ $2$ $2$ $2$ 00 00 00 $-2$
$\pi_{2}^{\prime}$ $2$ $2$ $2$ 00 00 00 $2$
$\pi_{2}^{\prime\prime}$ $2$ $2$ $2$ 00 $-2$ 00 00
$\pi_{2}^{\prime\prime\prime}$ $2$ $2$ $2$ 00 $2$ 00 00
@@
$\pi_{9}^{\prime\prime\prime}$ $9$ $9$ $1$ $-1$ $-1$ $1$ $-1$
@@ E.4 Example at $n=g=2$
in the notation of table 2. Thus we find that $\Theta(\pi_{1})=r_{1}+r_{15}^{\prime}+r_{24}$, confirming that $\pi_{1}\otimes r_{1}$ only appears once inside $\mathcal{X}$, and that $\Theta(\pi_{1})$ contains exactly $31+\min(n,g)=$ 3 terms in agreement with (122), where $r_{1}=\mathbf{1}$, $r_{24}=r_{[1],[1]}$ and $r_{15}^{\prime}=r_{\emptyset,[2]}$. In all the examples above, we had $21+\min(n,g)=$ 2, and there were two irreps in $\Theta(\pi_{1})$.
@@
even codewords $c$ of length $n$, i.e. vectors in ${\mathscr{D}}={\mathbb{F}}_{p}^{2n}$ satisfying (43). To derive that, we first consider the case of $c=(\vec{\alpha},\vec{\beta})$ with $\vec{\alpha}\neq 0$. There are $p^{n}-1$ such vectors. For each such vector $\vec{\alpha}$ we can have arbitrary vector $\vec{\beta}$ of length $n-1$. One of $n$ entries of $\vec{\beta}$ will be fixed to assure $\vec{\alpha}\cdot\vec{\beta}=0$ (mod) $p$. This entry could be not the last one, but should correspond to $\alpha_{i}\neq 0$. In the end, there are $p^{n-1}$ such vectors $\vec{\beta}$. Hence we have $(p^{n}-1)p^{n-1}$ pairs. Next, we consider zero vector $\vec{\alpha}$. Then any vector $\vec{\beta}$ works – there are $p^{n}$ such possibilities. We thus have ${\mathcal{N}}(p)=(p^{n}-1)p^{n-1}+p^{n}$.
@@
Here we use 00{0} to denote the zero element inside ${\mathscr{D}}$. A sum here is over all non-zero even codewords. It can be represented as a sum over all $p^{2n}$ vectors $c=(\vec{\alpha},\vec{\beta})\in{\mathscr{D}}={\mathbb{F}}_{p}^{2n}$ subject to the constraint (43), which can be imposed by an additional sum,
@@
If we instead sum $\psi_{r_{1}c,r_{2}c}$ for all even non-zero $c$ and arbitrary $0\leq r_{1},r_{2}<p$, we should additionally divide by $(p-1)$ because each non-zero $c_{1}=r_{1}c$ will appear $p-1$ times. Let us also note that in the sum $\psi_{r_{1}c,r_{2}c}$ over $c,r_{1},r_{2}$, the number of terms $\psi_{0,c}$ and $\psi_{c,0}$ will be $(p-1)$ for each non-zero even $c$. The number of such terms in each $\overline{W}$ should be $p^{n}-1$, hence coefficient (237) is correct for terms $\Psi_{{0},c}$ and $\Psi_{c,0}$ as well. We therefore find
@@
Here the sums over $r,r_{i},l$ are from 00 to $p-1$, and the sums over $\vec{\alpha}$, $\vec{\beta}$ etc are over ${\mathbb{F}}_{p}^{n}$.

### v2

@@ Appendix D Higher-genus modular transformations and MacWilliams identities
Just as the genus-1 enumerator polynomials of even self-dual codes are invariant under the $S$ and $T$ transformations (47) and (48), which correspond to the genus-$1$ modular group, the genus-$g$ enumerator polynomial are invariant under transformations of the genus-$g$ modular group $Sp(2g,{\mathbb{Z}})$. Here we give explicit form of the generators of $\mathrm{Sp}(2g,{\mathbb{F}}_{p})$ acting on genus-$g$ codeword blocks (Chern–Simons wavefunctions), or equivalently, variables entering enumerator polynomials. All relations here in fact generalize to groups defined over ${\mathbb{Z}}_{k}$ for general $k$.
@@
$\Psi_{c_{1},\ldots,c_{g}}(\Omega,\xi,\bar{\xi})=\frac{1}{\Phi(\Omega)}\sum_{v_{I}\in\Lambda_{c_{I}}}e^{i\pi\,v_{I}^{T}\Pi_{IJ}v_{J}+2\pi i(p^{I}_{L}\cdot\xi^{I}-p^{I}_{R}\cdot\bar{\xi}^{I})+\pi\,\Xi_{IJ}(\xi^{T}_{I}\xi_{J}+\bar{\xi}^{T}_{I}\bar{\xi}_{J})},$ (188)
@@
$\Psi_{c_{1},\ldots,c_{g}}(\Omega^{\prime},\xi^{\prime},\bar{\xi}^{\prime})$ $=$ $\Psi_{c_{1},\ldots,c_{g}}(\Omega,\xi,\bar{\xi})\,e^{\frac{i\pi}{p}(c_{I},c_{J})T_{IJ}}.$ (191)
The corresponding $Sp(2g,{\mathbb{F}}_{p})$ group element is $\left(\begin{smallmatrix}1&T\\ 0&1\end{smallmatrix}\right)$.
@@
$\Omega\rightarrow(\Omega^{\prime})_{ij}$ $=\left\{\begin{array}[]{cr}\Omega_{IJ}-\Omega_{IL}\Omega_{LJ}/\Omega_{LL},&I,J\neq L,\\ \Omega_{IL}/\Omega_{LL},&J=L,\\ \Omega_{LJ}/\Omega_{LL},&I=L,\\ -1/\Omega_{LL},&I=J=L,\end{array}\right$. (196)
$\xi\rightarrow\xi^{\prime}_{i}$ $=\left\{\begin{array}[]{cr}\xi_{I}-\xi_{L}\Omega_{LI}/\Omega_{LL},&I\neq L,\\ \xi_{L}/\Omega_{LL},&I=L,\end{array}\right$. (199)
$\bar{\xi}\rightarrow\bar{\xi}^{\prime}_{i}$ $=\left\{\begin{array}[]{cr}\bar{\xi}_{I}-\bar{\xi}_{L}{\bar{\Omega}}_{LI}/\bar{\Omega}_{LL},&I\neq L,\\ \bar{\xi}_{L}/\bar{\Omega}_{LL},&I=L,\end{array}\right$. (202)
@@
This transformation corresponds to the $Sp(2g)$ matrix
@@ Appendix E Explicit examples
In this appendix, we present additional examples that generalize the simplest non-trivial case of $[png]=[311]$ of section 3.3. We begin by presenting some commands in GAP GAP that are useful for working out the examples. Then we give complete examples of the theta correspondence for the cases $[png]=[312],[511],[321]$, and partial results for the case $[png]=[322]$.
@@ E.1 Method and manipulations in GAP
First we explain how to perform some of the computations behind the examples. In summary, we need to exhibit a finite-dimensional reducible representation of the group $G\times H$ on the space $\mathcal{X}$ of dimension $p^{2ng}$ whose basis are the vectors $\psi_{c_{1},\ldots,c_{g}}$. We take $G=O(n,n,{\mathbb{F}}_{p})=O^{+}(2n,{\mathbb{F}}_{p})$ (the meaning of ++ was explained in appendix B) and $H=Sp(2g,{\mathbb{F}}_{p})$. These groups can be generated in GAP by specifying the quadratic/bilinear form. Here we give the commands for $[png]=[311]$. The orthogonal group and its character table are generated by the following commands: [^38] In GAP, double semicolon suppresses the output. We found that for the case $p=5$, $n=2$, $g=1$ there are some issues in executing the code given here. Note also that in GAP, as in ATLAS Atlas1985 , the orthogonal group is denoted “GO” (general orthogonal).
@@ E.1.1 Example at genus $g=2$
The space $\mathcal{X}_{+}^{[312]}$ has dimension $3^{4}=81$, [^41] Comparing with (169) ($N=4$), it is a direct sum of two irreps of the larger symplectic group $Sp(8,{\mathbb{F}}_{3})$ with dimensions $40$ (charge-conjugation odd) and $41$ (charge-conjugation even) respectively. and following the steps above, we find that under the product group $O^{+}(2,{\mathbb{F}}_{3})\times Sp(4,{\mathbb{F}}_{3})$ it has a decomposition,
@@
$Sp_{4,3}$ $[1]$ $1]^{\prime} [$90][90] $[4320]$ $[6480]$ $[40]$ $\cdots$
$r_{1}$ 11 11 11 11 11 11
$r_{4}$ 44 4-4 0 0 0 $2\zeta-\zeta^{2}$
$r_{4}^{\prime}$ 44 4-4 0 0 0 $2\zeta^{2}-\zeta$
$r_{5}$ 55 55 3-3 11 1-1 $\zeta^{2}-2\zeta$
$r_{5}^{\prime}$ 55 55 3-3 11 1-1 $\zeta-2\zeta^{2}$
$r_{6}$ 66 66 2-2 1-1 0 3-3
$r_{10}$ $10$ $10$ 22 11 0 $2\zeta^{2}+5\zeta$
$r_{10}^{\prime}$ $10$ $10$ 22 11 0 $2\zeta+5\zeta^{2}$
$r_{15}$ $15$ $15$ 1-1 1-1 1-1 66
$r_{15}^{\prime}$ $15$ $15$ 77 0 11 3-3
$r_{20}$ 2020 20-20 0 0 0 7-7
$r_{20}^{\prime}$ 2020 2020 44 11 0 22
$r_{20}^{\prime\prime}$ 2020 20-20 0 0 0 $5\zeta-8\zeta^{2}$
$r_{20}^{\prime\prime\prime}$ 2020 20-20 0 0 0 $5\zeta^{2}-8\zeta$
$r_{20}^{\prime\prime\prime\prime}$ 2020 20-20 0 0 0 $\zeta-5\zeta^{2}$
$r_{20}^{\prime\prime\prime\prime\prime}$ 2020 20-20 0 0 0 $\zeta^{2}-5\zeta$
$r_{24}$ $24$ $24$ 88 0 0 66
$r_{30}$ $30$ $30$ 10-10 1-1 0 33
@@
Table 2: Truncated character table for the group $Sp(4,{\mathbb{F}}_{3})$ with $51840$ elements, displaying the two conjugacy classes with one element ($[1]$ and $1]^{\prime}$), three additional conjugacy classes whose size is unique among the conjugacy classes, and one additional conjugacy class necessary to distinguish the representations $r_{20}^{\prime\prime}$ and $r_{20}^{\prime\prime\prime}$. The whole character table contains $34$ conjugacy classes. $\zeta=e^{2\pi i/3}$.
@@
[…]
@@ E.2 Example for prime $p=5$
where the last term constitutes the charge-conjugation odd subspace of dimension $12$. Again, we see that $\pi_{1}\otimes r_{1}$ only appears once. The identification with bipartitions for unipotent irreps is the same as in the example $p=3$ in the main text, where now the dimension formula (118) shows that $\dim r_{\emptyset,[1]}=5$.
$O_{2,5}^{+}$ $[1]$ $[2]$ $^{\prime} [2]′′[2]^{\prime\prime}$ $1]^{\prime}$
$\pi_{1}$ 11 11 11 11 11
$\pi_{1}^{\prime}$ 11 1-1 1-1 11 11
$\pi_{1}^{\prime\prime}$ 11 1-1 11 1-1 11
$\pi_{1}^{\prime\prime\prime}$ 11 11 1-1 1-1 11
$\pi_{2}$ 22 0 0 0 2-2
@@
$SL_{2,5}$ $[1]$ $[12]$ $2]^{\prime} [1][1$$^{\prime}$ $^{\prime\prime} [12]′′′[12]^{\prime\prime\prime}$ $[20]$ $20]^{\prime} [$30][30]
@@
$r_{2}$ 2 $\xi^{4}-\xi$ $\xi^{3}-\xi^{2}$ 2-2 $\xi^{4}+\xi$ $\xi^{3}+\xi^{2}$ 1-1 11 0
$r_{2^{\prime}}$ 22 $\xi^{3}-\xi^{2}$ $\xi^{4}-\xi$ 2-2 $\xi^{3}+\xi^{2}$ $\xi^{4}+\xi$ 1-1 11 0
$r_{3}$ 33 $\xi^{3}-\xi^{2}$ $\xi^{4}-\xi$ 33 $\xi^{3}-\xi^{2}$ $\xi^{4}-\xi$ 0 0 1-1
$r_{3^{\prime}}$ 33 $\xi^{4}-\xi$ $\xi^{3}-\xi^{2}$ 33 $\xi^{4}-\xi$ $\xi^{3}-\xi^{2}$ 0 0 1-1
$r_{4}$ 44 1-1 1-1 44 1-1 1-1 11 11 0
$r_{4}^{\prime}$ 44 11 11 4-4 1-1 1-1 11 1-1 0
$r_{5}$ 55 0 0 55 0 0 1-1 1-1 11
$r_{6}$ 66 1-1 1-1 6-6 11 11 0 0 0
@@ E.3 Example at length $n=2$
$O_{4,3}^{+}$ $[1]$ $1]^{\prime} [$18][18] $[144]$ $[72]$ $^{\prime} [72]′′[72]^{\prime\prime}$ $\cdots$
$\pi_{1}$ 11 11 11 11 11 11 11
@@
$\pi_{2}$ 22 22 22 0 0 0 2-2
$\pi_{2}^{\prime}$ 22 22 22 0 0 0 22
$\pi_{2}^{\prime\prime}$ 22 22 22 0 2-2 0 0
$\pi_{2}^{\prime\prime\prime}$ 22 22 22 0 22 0 0
@@
$\pi_{9}^{\prime\prime\prime}$ 99 99 11 1-1 1-1 11 1-1
@@ E.4 Example at $n=g=2$
in the notation of table 2. Thus we find that $\Theta(\pi_{1})=r_{1}+r_{15}^{\prime}+r_{24}$, confirming that $\pi_{1}\otimes r_{1}$ only appears once inside $\mathcal{X}$, and that $\Theta(\pi_{1})$ contains exactly $1+\min(n,g)=3$ terms in agreement with (122), where $r_{1}=\mathbf{1}$, $r_{24}=r_{[1],[1]}$ and $r_{15}^{\prime}=r_{\emptyset,[2]}$. In all the previous examples, we had $1+\min(n,g)=2$, and there were two irreps in $\Theta(\pi_{1})$.
@@
even codewords $c$ of length $n$, i.e. vectors in ${\mathscr{D}}={\mathbb{F}}_{p}^{2n}$ satisfying (43). To derive that, we first consider the case of $c=(\vec{\alpha},\vec{\beta})$ with $\vec{\alpha}\neq 0$. There are $p^{n}-1$ such vectors. For each such vector $\vec{\alpha}$ we can have arbitrary vector $\vec{\beta}$ of length n1n-1. One of $n$ entries of $\vec{\beta}$ will be fixed to assure $\vec{\alpha}\cdot\vec{\beta}=0$ (mod) $p$. This entry could be not the last one, but should correspond to $\alpha_{i}\neq 0$. In the end, there are $p^{n-1}$ such vectors $\vec{\beta}$. Hence we have $(p^{n}-1)p^{n-1}$ pairs. Next, we consider the case of zero vector $\vec{\alpha}$. Then any vector $\vec{\beta}$ works – there are $p^{n}$ such possibilities. We thus have ${\mathcal{N}}(p)=(p^{n}-1)p^{n-1}+p^{n}$.
@@
Here we use 0{0} to denote the zero element inside $\mathscr{D}$}. A sum here is over all non-zero even codewords. It can be represented as a sum over all $p^{2n}$ vectors $c=(\vec{\alpha},\vec{\beta})\in{\mathscr{D}}={\mathbb{F}}_{p}^{2n}$ subject to the constraint (43), which can be imposed by an additional sum,
@@
If we instead sum $\psi_{r_{1}c,r_{2}c}$ for all even non-zero $c$ and arbitrary $0\leq r_{1},r_{2}<p$, we should additionally divide by (p1)(p-1) because each non-zero $c_{1}=r_{1}c$ will appear p1p-1 times. Let us also note that in the sum $\psi_{r_{1}c,r_{2}c}$ over $c,r_{1},r_{2}$, the number of terms $\psi_{0,c}$ and $\psi_{c,0}$ will be (p1)(p-1) for each non-zero even $c$. The number of such terms in each $\overline{W}$ should be $p^{n}-1$, hence coefficient (237) is correct for terms $\Psi_{{0},c}$ and $\Psi_{c,0}$ as well. We therefore find
@@
Here the sums over $r,r_{i},l$ are from 0 to p1p-1, and the sums over $\vec{\alpha}$, $\vec{\beta}$ etc are over ${\mathbb{F}}_{p}^{n}$.